# V-LASIK: Consistent Glasses-Removal from Videos Using Synthetic Data

### v1

@@ Abstract.
Diffusion-based generative models have recently shown remarkable image and video editing capabilities. However, local video editing, particularly removal of small attributes like glasses, remains a challenge. Existing methods either alter the videos excessively, generate unrealistic artifacts, or fail to perform the requested edit consistently throughout the video.
In this work, we focus on consistent and identity-preserving removal of glasses in videos, using it as a case study for consistent local attribute removal in videos. Due to the lack of paired data, we adopt a weakly supervised approach and generate synthetic imperfect data, using an adjusted pretrained diffusion model. We show that despite data imperfection, by learning from our generated data and leveraging the prior of pretrained diffusion models, our model is able to perform the desired edit consistently while preserving the original video content. Furthermore, we exemplify the generalization ability of our method to other local video editing tasks by applying it successfully to facial sticker-removal. Our approach demonstrates significant improvement over existing methods, showcasing the potential of leveraging synthetic data and strong video priors for local video editing tasks.
@@ 1. Introduction
Recent advances in diffusion-based generative models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have demonstrated impressive capabilities in image and video editing (Rombach et al., 2022; Zhang et al., 2023a; Geyer et al., 2023; Yang et al., 2023; Zhang et al., 2023b; Kara et al., 2023; Bar-Tal et al., 2024). While local editing, such as removing and adding attributes without changing the rest of the content, mainly works for images, local video editing remains a challenge. Video frames often contain motion blur and challenging poses that are less common in images. Moreover, videos of people pose additional challenges, as humans are highly sensitive to subtle unrealistic “uncanny” artifacts, and our visual system is more sensitive to motion inconsistencies, so the quality bar is higher. Existing image editing and inpainting methods (Nitzan et al., 2024; Zhang et al., 2023a; Tsaban and Passos, 2023; Brooks et al., 2023; Couairon et al., 2023; Meng et al., 2021) are trained on clean images, that do not contain the aforementioned problems, and therefore do not perform well over video frames; They typically either change the original frames too much, or do not perform the requested change completely, as illustrated in Fig. 2.
Video inpainting methods (Zhou et al., 2023; Yu et al., 2023a; Li et al., 2022b; Zhang et al., 2022) typically remove objects by deleting and filling regions of the video, often disregarding the information originally present in those regions. When it comes to generating new video content behind an object that never moves, these methods tend to generate a low detailed background, or one with clear unrealistic artifacts. As a representing use-case, we choose to tackle glasses removal, as it is a particularly challenging example; where some content is never or only partially revealed (e.g. eyes, eyebrows), and is changed across the video (e.g. eyelids change position across frames). Since eyes contribute greatly to identity perception, a consistent and highly realistic generation is required for an identity to be preserved. However, current video inpainting works typically generate smooth results when using small masks, that do not fit the facial case well (see ProPainter (Zhou et al., 2023) and FGT (Zhang et al., 2022) results in Fig. 5).
Mask-free video editing methods (Geyer et al., 2023; Zhang et al., 2023b; Khachatryan et al., 2023; Kara et al., 2023; Cong et al., 2024) mainly focus on changing the style of the whole video, or changing the appearance of large parts of it, and not on local changes that should adhere to the original context. Other works (Kasten et al., 2021; Bar-Tal et al., 2022) learn an atlas for each new video, leading to long inference time. Furthermore, as mentioned in Text2Live (Bar-Tal et al., 2022), the goal of these works is to manipulate the appearance of existing objects, not to change the composition of the elements in the scene. Hence, their method is not capable of successfully removing attributes such as glasses from faces.
In this work, we focus on removing glasses from videos as a case study for local video editing. We introduce a new method that performs this task consistently while preserving the original identity. To the best of our knowledge, no paired data exists for this task; Therefore, we cannot finetune a model directly to solve it. To tackle this problem, we generate imperfect synthetic data, and use it to train a model in a weakly supervised manner. We show that by using an adjusted pretrained image inpainting model, we can generate imperfect data pairs, and learn from them. Our method is closest to video inpainting, in the sense that it removes something from the video and replaces it with different content. However, similarly to recent video editing works, the input to our model is a mask-free image, allowing the model to use facial details that are occluded for inpainting models. As glasses often contain colored lens and reflections that change across frames, which affect the way the eye behind them looks, they are a challenging attribute to remove from faces. Hence, we incorporate cross-frame attention layers in our data generation model, which allows aggregating information from different frames. Still, as presented in Fig. 2, the generated data (‘Synth data’) is imperfect and contains many artifacts, such as not always preserving the eyelid position and other features from the original frame. To overcome these challenges, we finetune a diffusion based image-to-image model over this data. This way we can leverage the prior of the pretrained model, while enhancing its capabilities with our new data. Additionally, the non-masked input allows the model to exploit the fact that it is exposed to useful information within the masked region to achieve the requested edit, while preserving the rest of the content from the original frames. Despite the imperfections in our synthetic data, a model with a strong prior can leverage its prior knowledge to outperform the training data and produce high-quality results. Finally, to get temporally consistent results, we combine our trained model with a motion prior, and achieve state-of-the-art results on video glasses-removal, surpassing all current video editing and inpainting methods. To demonstrate that our method can be generalized to other local video editing tasks, we successfully apply it to the task of removing stickers from faces. We release code and both the stickers and the glasses removal datasets for further research.
@@ 2.1. Image editing
Since the advent of diffusion models for image generation and editing (Ho et al., 2020), real image editing has rapidly advanced and showed remarkable results (Brooks et al., 2023; Couairon et al., 2023; Tsaban and Passos, 2023; Hertz et al., 2022; Tumanyan et al., 2023; Zhang et al., 2023a; Meng et al., 2021), including local editing such as removing glasses from images. Moreover, image inpainting methods (Nitzan et al., 2024; Rombach et al., 2022; Zhang et al., 2023a; Yildirim et al., 2023), have also shown impressive results, realistically replacing the content behind a given mask. Additionally, prior work (Lyu et al., 2022; Lee and Lai, 2020) tried to tackle the task of removing glasses from images directly. Indeed, these works perform well over images. However, as shown in Fig. 2 they fail when applied to video frames of people, that do not always look directly at the camera, and constantly move, causing motion blur and other artifacts. Moreover, when applied frame-by-frame, the generated content differs between frames, resulting in temporal inconsistencies.
@@ 2.2. Video editing
Recently, video editing has also developed greatly with video-to-video translation methods (Geyer et al., 2023; Khachatryan et al., 2023; Cong et al., 2024; Kara et al., 2023; Qi et al., 2023; Wu et al., 2023; Zhang et al., 2023b) that focus on transforming the entire frame into a different style, while trying to preserve temporal consistency in the generated videos. Some of them, such as RAVE (Kara et al., 2023), or TokenFlow (Geyer et al., 2023) with SDEdit (Meng et al., 2021) also perform local editing. However, in the case of glasses-removal, where a person moves across the video and the requested edit is small and delicate, many artifacts such as face deformations and inconsistencies are generated. Examples are presented in Fig. 5 and in video results in the supplementary material. Another line of works is atlas-based video editing (Kasten et al., 2021; Lu et al., 2021; Bar-Tal et al., 2022; Suhail et al., 2023). These methods learn an atlas for each video, to apply changes to either the background or foreground of the video. The initial atlas reconstruction requires excessive computational resources and long running times. Moreover, as these methods were designed to change the appearance of existing objects, they are not meant for adding or removing attributes that were not part of the original video, such as glasses, and do not perform these changes well. Additionally, while theoretically allowing to remove layers from a video, these methods are limited to relatively simple motion and to static backgrounds.
@@ 2.3. Video inpainting
Video inpainting is the task of consistently filling-in new content behind a given mask throughout a video (Yu et al., 2023a; Zhang et al., 2022; Zhou et al., 2023; Gu et al., 2023; Li et al., 2022b). Such methods work well when the object moves throughout the video, and the model is able to fill-in the background from other video frames, where the background is visible. However, they struggle with filling-in completely occluded areas, especially when they are part of the foreground, such as faces. Particularly, areas behind glasses are only partially revealed by head motion, and sometimes glasses cause optical deformations that hide the background. Slight changes to face features in the “background” of glasses may yield unrealistic results, or alter the face identity, as shown in the ProPainter (Zhou et al., 2023) and FGT (Zhang et al., 2022) results in Fig. 5. While such changes to smooth backgrounds may not significantly affect the perceptual quality of the results, in foreground areas and particularly in faces they are detrimental.
@@ 2.4. Learning from synthesized data
Deep learning models require large amount of data to learn from, however paired data is not always available for the task at hand. Hence, several works (Li et al., 2022a; Peebles et al., 2022; Ravuri and Vinyals, 2019; Brooks et al., 2023; Lyu et al., 2022) use generative models for data generation to train models. These works usually rely on the high quality of the generated data, and learn solely from it. In this work, we accept that generated data is usually imperfect, and is insufficient to achieve the required result on its own, thus we take advantage of the strong priors of pretrained models and use our generated data for fine-tuning. This way, the model learns to generate results that are superior to the data it was trained on, while performing the relevant task.
Lyu et al. (Lyu et al., 2022) specifically explores synthetic data for the task of glasses-removal from images. However, their data acquisition process requires face scanning and 3D data, which is hard to acquire, unlike our method which does not require any special equipment or effort.
@@ 3.2. Model fine tuning
After the data generation stage, we obtain pairs of frames with and without glasses. However, our generated frames are not temporally consistent, and also contain per-frame artifacts in many cases, as featured in Figs. 2 and 4. For example, the eyelid positions, i.e. closed or open, sometimes change after the inpainting process. Despite that, the generated data is good enough for finetuning a pretrained image-to-image diffusion model for the task of removing glasses from faces. These models have strong prior for generating realistic looking images that are similar to the original ones and preserve their small and delicate details. Therefore, after finetuning, the model learns the task of removing glasses from our data, while preserving fine details such as eye color and eyelid positions.
@@ 4. Implementation Details
General. We use Stable Diffusion 1.5 (rombach2021highresolution) as our backbone in all the method steps, as at the time of development there was no compatible motion module for SDXL.
@@ 5.3. Quantitative evaluation
Identity preservation: to test identity preservation, we use an ID score ($ID$ in Tab. 1), which is defined by the average cosine similarity between the face embeddings of the video frames, generated by the face recognition model Arcface (Deng et al., 2019). We note that neither of these metrics is complete on its own, as an unchanged video would get a perfect $ID$ score, and a random video without glasses would get a very high $\Delta G$ score. For example, as Text2Video-Zero (Khachatryan et al., 2023) does not remove the glasses from most videos, and does not change the videos by much, it achieves a very high $ID$ score. For this reason, we follow prior work (Kara et al., 2023; cong2023flatten) and also look at $ID\cdot\Delta G$, which quantifies the trade-off between removing the glasses from the video, and remaining faithful to the identity of the person in the original video. When looking at $ID\cdot\Delta G$, our model achieves the best results, together with ProPainter (Zhou et al., 2023). However, as mentioned, the results of our user-study in Fig. 7 show that our model outperforms ProPainter and all other methods both in terms of glasses-removal, and in terms of realism and identity preservation.
@@ 7. Societal impact
While the goal of this work is to allow editing of owned or licensed videos only, we acknowledge the fact it may be misused to altering videos without consent, and contribute to the spread of misinformation. We condemn such usage, and we are actively working on systems that detect synthetic and edited media.
@@ 9. Acknowledgements
This project was funded in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreement No. 802774 (iEXTRACT), and by ISF grant number 1337/22.

### v2

@@ Abstract.
Recently, diffusion-based generative models showed remarkable image and video editing capabilities. However, local video editing, particularly removal of small attributes like glasses, remains a challenge. Existing methods either alter the videos excessively, generate unrealistic artifacts, or fail to perform the requested edit consistently. In this work, we focus on consistent and identity-preserving glasses-removal in videos, as a case study for video local attribute removal. We demonstrate the generalizability of our method by applying it to facial sticker removal from videos. Due to the lack of paired data, we adopt a weakly supervised approach and generate synthetic imperfect data, using an adjusted pretrained diffusion model. We show that despite data imperfection, by learning from our generated data and leveraging the prior of pretrained models, our model is able to perform the desired edit consistently while preserving the original video content. Furthermore, we suggest a new normalization method, Inside-Out Normalization, which aligns colors in a filled region with colors outside that region. Our approach offers significant improvement over existing methods, showcasing the potential of leveraging synthetic data and strong priors for local video editing tasks.
@@ 1. Introduction
Recent advances in diffusion-based generative models Ho et al. (2020); Sohl-Dickstein et al. (2015) have demonstrated impressive capabilities in image and video editing (Rombach et al., 2022; Zhang et al., 2023a; Geyer et al., 2023; Yang et al., 2023; Zhang et al., 2023b; Kara et al., 2023; Bar-Tal et al., 2024). While local editing, such as removing attributes without changing the rest of the content, mainly works for images, for videos it remains a challenge. Video frames often contain motion blur and challenging poses that are less common in images. Moreover, videos of people pose additional challenges, as humans are highly sensitive to subtle unrealistic “uncanny” artifacts, and our visual system is more sensitive to motion inconsistencies, which leads to a higher quality bar. Existing image editing and inpainting methods Meng et al. (2021); Zhang et al. (2023a); Tsaban and Passos (2023); Brooks et al. (2023); Couairon et al. (2023); Nitzan et al. (2024) are trained on images that do not contain the aforementioned problems, and therefore do not perform well over video frames; They either change the frames too much, or do not perform the requested edit completely, as illustrated in Fig. 2. Video inpainting methods Zhou et al. (2023); Yu et al. (2023a); Li et al. (2022b); Zhang et al. (2022) typically remove objects by deleting and filling regions of the video, often disregarding the information originally present in those regions. When trying to generate new video content behind an object that never moves, these methods tend to generate a low detailed background, or one with clear unrealistic artifacts. As a representing use-case for local attribute removal, we choose to tackle glasses removal, as it is a particularly challenging example; where some content is never or only partially revealed (e.g. eyes, eyebrows), and is changed across the video (e.g. eyelids change position). Since eyes contribute greatly to identity perception, a consistent and highly realistic generation is required for an identity to be preserved. However, current video inpainting works typically generate smooth results when using small masks, that do not fit the facial case well (see ProPainter Zhou et al. (2023), FGT Zhang et al. (2022) results in Fig. 5). Mask-free video editing methods Geyer et al. (2023); Zhang et al. (2023b); Khachatryan et al. (2023); Kara et al. (2023); Cong et al. (2024) mainly focus on changing the style of the whole video or large parts of it, and not on local changes that should adhere to the original context. Other works Kasten et al. (2021); Bar-Tal et al. (2022) learn an atlas for each new video, leading to long inference times. Furthermore, the goal of these works is to manipulate the appearance of existing objects, not to change the composition of the elements in the scene (Bar-Tal et al., 2022). Hence, their method is not capable of successfully removing attributes such as glasses from faces.
@@
In this work, we introduce a new method that consistently removes glasses from videos while preserving the content of the original video. To the best of our knowledge, no paired data exists for this task; Therefore, we cannot finetune a model directly to solve it. To tackle this problem, we generate imperfect synthetic data, and use it to train a model in a weakly supervised manner. We show that by using an adjusted pretrained image inpainting model, we can generate imperfect data pairs, and learn from them. Our method is closest to video inpainting in the sense that it removes something from the video and replaces it with different content. However, similarly to recent video editing works, the input to our model is a mask-free image, exposing it to facial details that are occluded for inpainting models. As glasses often contain colored lens and reflections that change across frames, which affect the way the eye behind them looks, they are a challenging attribute to remove from faces. Hence, we incorporate cross-frame attention layers in our data generation model, which allows aggregating information from different frames. Still, as presented in Fig. 2, the generated data (‘Synth data’) is imperfect and contains many artifacts, such as not always preserving the eyelids positions from the original frame. To overcome these challenges, we finetune a pretrained diffusion based image-to-image model over this data. This way, we can leverage its prior, while enhancing its capabilities with our new data. Despite imperfections in our synthetic data, the model can leverage its prior knowledge to outperform the training data and produce high-quality results. For temporally consistent results, we combine our trained model with a motion prior. Finally, for smooth and content-faithful edits, we suggest a novel normalization method, Inside-Out Normalization (ION) that allows matching the colors of a masked region with the area outside the mask. We achieve state-of-the-art results on video glasses-removal, surpassing all current video editing and inpainting methods. To demonstrate that our method can be generalized to other local video editing tasks, we successfully apply it to the task of removing stickers from faces.
@@ 2.1. Image editing
Since the advent of diffusion models for image generation and editing Ho et al. (2020), real image editing has rapidly advanced and showed remarkable results Brooks et al. (2023); Couairon et al. (2023); Tsaban and Passos (2023); Hertz et al. (2022); Tumanyan et al. (2023); Zhang et al. (2023a); Meng et al. (2021), including local editing such as glasses-removal. Image inpainting methods Nitzan et al. (2024); Rombach et al. (2022); Zhang et al. (2023a); Yildirim et al. (2023), also showed impressive results, realistically replacing content behind a given mask. Moreover, prior work Lyu et al. (2022); Lee and Lai (2020) tackled the task of image glasses-removal directly. These works perform well over images. However, as shown in Fig. 2 they fail when applied to video frames of people, that do not always look directly at the camera, and constantly move, causing motion blur and other artifacts. Moreover, when applied frame-by-frame, the generated content differs between frames, resulting in temporal inconsistencies. Prior to diffusion models, Generative Adversarial Networks (GANs) have been widely used for facial image editing (Karras et al., 2019, 2020; Yao et al., 2021; Patashnik et al., 2021; Richardson et al., 2021). However, In addition to temporal consistency issues, they operate on aligned crops of faces, thus when applied to videos, faces must be stitched back to the original frames which is challenging and often yields unsatisfactory results.
@@ 2.2. Video editing
Recently, video editing has developed greatly with video-to-video translation methods Geyer et al. (2023); Khachatryan et al. (2023); Cong et al. (2024); Kara et al. (2023); Qi et al. (2023); Wu et al. (2023); Zhang et al. (2023b) that focus on transforming the entire frame into a different style, while trying to preserve temporal consistency in the generated videos. Some of them, such as RAVE Kara et al. (2023), or TokenFlow Geyer et al. (2023) with SDEdit Meng et al. (2021) also perform local editing. However, in the case of glasses-removal, where a person moves across the video and the requested edit is small and delicate, many artifacts such as face deformations and inconsistencies are generated. Examples are presented in Fig. 5 and in video results in the supplementary material. Earlier work used GANs to perform local editing of faces (Tzaban et al., 2022; Xu et al., 2022, 2023). Another line of works is atlas-based video editing (Kasten et al., 2021; Lu et al., 2021; Bar-Tal et al., 2022; Suhail et al., 2023). These methods learn an atlas for each video, to apply changes to either the background or foreground of the video. The initial atlas reconstruction requires excessive computational resources and long running times. Moreover, as these methods were designed to change the appearance of existing objects, they are not meant for adding or removing attributes that were not part of the original video, such as glasses, and do not perform these changes well.
@@ 2.3. Video inpainting
Video inpainting is the task of consistently filling-in new content behind a given mask throughout a video (Yu et al., 2023a; Zhang et al., 2022; Zhou et al., 2023; Gu et al., 2023; Li et al., 2022b). Such methods work well when the object moves throughout the video, and the model is able to fill-in the background from other video frames, where the background is visible. However, they struggle with filling-in completely occluded areas, especially when they are part of the foreground, such as faces. Particularly, areas behind glasses are only partially revealed by head motion, and sometimes glasses cause optical deformations that hide the background. Slight changes to face features in the “background” of glasses may yield unrealistic results, or alter face identity, as shown in the ProPainter Zhou et al. (2023) and FGT Zhang et al. (2022) results in Fig. 5. While such changes to smooth backgrounds may not affect the quality of the results significantly, in foreground areas, particularly in faces, they are detrimental.
@@ 2.4. Learning from synthesized data
Deep learning models require large amount of data to learn from, however paired data is not always available for the task at hand. Hence, several works Li et al. (2022a); Peebles et al. (2022); Ravuri and Vinyals (2019); Brooks et al. (2023); Lyu et al. (2022) used generative models for data generation to train models. These works usually rely on the high quality of the generated data, and learn solely from it. In this work, we accept that generated data is usually imperfect, and is insufficient to achieve the required result on its own, thus we take advantage of strong priors of pretrained models and use our generated data for fine-tuning. This way, the model learns to generate results that are superior to the data it was trained on, while performing the relevant task. Lyu et al. (2022) specifically explores synthetic data for the task of glasses-removal from images. However, their data acquisition process requires face scanning and 3D data, which is hard to acquire, unlike our method which does not require any special equipment or effort.
@@ 3.2. Model fine tuning
After the data generation stage, we obtain pairs of frames with and without glasses. However, our generated frames are not temporally consistent, and also contain per-frame artifacts in many cases, as featured in Figs. 2 and 4. For example, the eyelid positions, i.e. closed or open, sometimes change after the inpainting process. Despite that, the generated data is good enough for finetuning a pretrained image-to-image diffusion model for the task of removing glasses from faces.
These models have strong prior for generating realistic looking images that are similar to the original ones and preserve their small and delicate details. Therefore, after finetuning, the model learns the task of removing glasses from our data, while preserving fine details such as eye color and eyelid positions.
@@ 4. Implementation Details
General. We use Stable Diffusion 1.5 (Rombach et al., 2022) as our backbone in all the method steps, as at the time of development there was no compatible motion module for SDXL.
@@ 5.3. Quantitative evaluation
Identity preservation: to test identity preservation, we use an ID score ($ID$ in Tab. 1), which is defined by the average cosine similarity between the face embeddings of the video frames, generated by the face recognition model Arcface (Deng et al., 2019). We note that neither of these metrics is complete on its own, as an unchanged video would get a perfect $ID$ score, and a random video without glasses would get a very high $\Delta G$ score. For example, as Text2Video-Zero (Khachatryan et al., 2023) does not remove the glasses from most videos, and does not change the videos by much, it achieves a very high $ID$ score. For this reason, we follow prior work (Kara et al., 2023; Cong et al., 2024) and also look at $ID\cdot\Delta G$, which quantifies the trade-off between removing the glasses from the video, and remaining faithful to the identity of the person in the original video. When looking at $ID\cdot\Delta G$, our model achieves the best results, together with ProPainter (Zhou et al., 2023). However, as mentioned, the results of our user-study in Fig. 7 show that our model outperforms ProPainter and all other methods both in terms of glasses-removal, and in terms of realism and identity preservation.
@@ 7. Societal impact
While the goal of this work is to allow editing of owned or licensed videos only, we acknowledge the fact it may be misused to altering videos without consent, and contribute to the spread of misinformation. We condemn such usage, and we are actively working on systems that detect synthetic and edited media (Agarwal et al., 2020; Knafo, 2022; Sinitsa and Fried, 2024).
@@ 9. Acknowledgements
We thank Yotam Nitzan for fruitful discussions. This project was funded in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreement No. 802774 (iEXTRACT), and by ISF grant numbers 1574/21, 1337/22.